OpenStudy (anonymous):

What is the half life of a radioative substance that decays by 6% in 7 years?

OpenStudy (anonymous):

half life is the period of time it takes to decay 50% of its original value. so what ration of 50% is 6% and apply that ratio to 7

OpenStudy (anonymous):

ummmm 50/6?

OpenStudy (anonymous):

So I make it about 58.3333

OpenStudy (anonymous):

yep 50/6 * 7

OpenStudy (anonymous):

Ohh! Thanks!

OpenStudy (anonymous):

hmm that is not what i got. let me try again.

OpenStudy (anonymous):

i got 78.4

OpenStudy (anonymous):

i think this problem is a little more complicated that ratios. the general formula for decay is \[A=A_0e^{rt}\] where r is the rate of decay (r will be negative) t is time, \[A_0\] is the initial amount and \[A\] is what you have after t years. we do not know r, but we know that when t = 7 you have decreased by 6%. decreasing by 6% means you have 94% =.94 of the original amount so set \[.94=e^{7r}\] and solve for r \[7r = ln(.94)\] \[r=\frac{ln.94}{7}=-.008839\] rounded now half life means you have half the original amount so set \[\frac{1}{2}=e^{-.00839t}\] or better still (to avoid rounding error) \[\frac{1}{2}=e^{\frac{ln(.94)}{7}t}\] and solve for t \[ln(\frac{1}{2})=\frac{ln(.94)t}{7}\] \[t=\frac{7ln(\frac{1}{2})}{ln(.94)}=78.4\] rounded

OpenStudy (anonymous):

i could be wrong of course, but i am fairly sure this is correct.

OpenStudy (anonymous):

ah yes what was I thinking. but isn't decay \[e^{-rt}\] and not \[e^{rt}\] because that's growth not decay so taking the log of both sides should give \[-7r=\ln(0.94)\] the rest follows?

OpenStudy (anonymous):

I agree with the answer given. here is a doc to clarify the above comment. sorry for the misunderstanding

OpenStudy (anonymous):

you can write \[A=A_0e^{-rt}\] it is just a matter of preference. my 'r' was negative. if you write it your way r is positive. it makes no difference.

OpenStudy (anonymous):

there is another way to solve this problem that amounts to the same thing. since you know it decreases by 6% in 7 years you could write the formula is \[A=A_0(.94)^{\frac{t}{7}}\] then set \[\frac{1}{2}=(.94)^{\frac{t}{7}}\] and solve for t, via \[ln(\frac{1}{2})=ln((.94)^{\frac{t}{7}})=\frac{t}{7}ln(.94)\] and still get \[t=\frac{7ln(\frac{1}{2})}{ln(.94)}\]

OpenStudy (anonymous):

Nice :D